ne : × | G In a large data set the 40th : × | a percentile to Data Value (M. s/95299/quizzes/117820/take/questions/1966887 ted: Jan 30 at 2:59pm uiz Instructions Χ 「¥ Flarworld Question 4 4 Given a sample: 1.6, 5.2,2.8,3.7,4.0 Whats the sample mean? 3.46 Whats the sample standard deviation? 1.35 (Round to two decimals) Whats the z-score for 2.8? (Round to two decimals) 0.2 Whats the percentile rank (an integer) for 2.8? Previous Subr No new data to save. Last checked at 5:35pm
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

Sample mean = (1.6+5.2+2.8+3.7+4.0)/5 = 3.46

Sample std deviation = sqrt(((1.6-3.46)^2 + ... + (4-3.46)^2 / 5-1) )= 1.35

Z score of 2.8 is (x-mu)/sigma = (2.8-3.46)/1.35 = -0.49

Percentile rank of 2.8 = 20 percentile ( why? out of the 5 numbers it is the 2nd highest number, and therefore will be the upper cutoff for the 20th percentile)